Use the techniques of differentiation to find dy over dx
(i) 3e to the power x = xy+xsquared +ysquared
(ii) y= 2xcubed over (xsquared - 4)squared
1
Expert's answer
2020-06-08T18:14:42-0400

Hello!

(i) We have,

  • 3ex = xy + x2 + y2

Differentiating both sides with respect to x,we get

  • 3ex = (x*dy/dx + y) + 2x +2y*dy/dx ..... (Applying chain rule for xy and y2)
  • 3ex = x*dy/dx + 2y*dy/dx +2x + y
  • 3ex - 2x - y = (dy/dx)*(x+2y)
  • dy/dx = (3ex -2x - y)/(x + 2y) ....(answer)

(ii) We have,

  • y = 2x3/(x2 - 4)2 ..... (i)

The question is in y=u/v form and we will use the formula(dy/dx = (v*du/dx - u*dv/dx)/v2)

So here u = 2x3 and v = (x2 - 4)2

Therefore, du/dx = 6x2 and dv/dx = 2(x2 - 4)*2x = 4x(x2 - 4) ....(applying chain rule)

Differentiating equation (i) with respect to x

  • dy/dx = (v*du/dx - u*dv/dx)/v2

Putting values of u,v,du/dx,dv/dx, we get

  • dy/dx = (x2 - 4)2*6x2 - 2x3*4x(x2 - 4)/(x2 - 4)4
  • dy/dx = (x2 - 4)2*6x2 - 8x4(x2 - 4)/(x2 - 4)4
  • dy/dx = x2(x2 - 4)*(6(x2 - 4) - 8x2)/(x2 - 4)4
  • dy/dx = x2*(6x2 - 24 - 8x2)/(x2 - 4)3
  • dy/dx = x2*( -24 - 2x2)/(x2 - 4)3
  • dy/dx = -2x2*( x2 + 24)/(x2 - 4)3

Hence, dy/dx = -2x2*( x2 + 24)/(x2 - 4)3 ......(answer)